# Cybersecurity of Quantum Key Distribution Implementations

### v2

@@ 6.3 Examples
6.3.1 Quantum Side-Channel Attacks
The Large Pulse attack [17]:
@@ 6.3.1 Quantum Side-Channel Attacks
The Injection-Locking attack [22]:
@@
The Bright Illumination attack [21]:
@@
6.3.2 Quantum State-Channel Attacks
The photon-number splitting attack [16]:
@@ 6.3.2 Quantum State-Channel Attacks
The Time-Shift attack [23]:
@@
Trojan-Pony Attack [55]:
@@
Imperfect Faraday Mirror Attack [68]:
@@
6.4 Application to Security Analyses
As shown in the subsection above, a wide variety of side channels can be used to attack QKD implementations. For an implementation to be proven secure, all these side channels must be considered in the theoretical model of the implementation, though we are not aware of security proofs that use this full modeling. Alternatively, a separate engineering analysis can examine the physical QKD device and its potential side channels and ensure that these side channels cannot be exploited by Eve.
7 Putting It All Together: Analysis of the Bright Illumination Attack
@@
7.1 The Bright Illumination Attack
The “Bright Illumination” attack [31, 21, 69, 56] is an implementation attack that changes the behavior of QKD receivers by exposing them to high-intensity light. This attack was implemented against a variety of QKD implementations, including implementations from recent years [70, 71], using various photodetectors and implementing various protocols [72, 73]. In this section we analyze the specific attack and implementation shown in [56]; however, our analysis can be applied with minor modifications to other Bright Illumination attacks against various QKD implementations.
@@ 7.1 The Bright Illumination Attack
In the QKD system targeted by [56], the implemented protocol is polarization-based BB84, and measurement is performed using a passive basis choice and four detectors, as shown in Figure 1 (a).
Figure 1: (a) Structure of the QKD receiver targeted in the Bright Illumination attack. PIBS: polarization-independent beam splitter. PBS: polarizing beam splitter. PR: polarization rotator. APD: avalanche photodiode. (b) Propagation of Eve’s high-intensity pulse through the QKD receiver.
@@
APDs have two possible modes of operation, called the “Geiger mode” and “linear mode” [31]. In Geiger mode, an APD operates as a single-photon detector: it generates a “click” (a macroscopic event) when it receives one or more photons. However, in linear mode, the APD “clicks” only if the incoming pulse’s intensity is above a certain threshold: if the pulse intensity is below that threshold, no effect is registered in the APD (there is no click). Thus, an APD in linear mode is “blind” to low-intensity pulses, including single photons.
In QKD setups, APDs operate in Geiger mode as single-photon detectors. However, an APD can be forcibly reconfigured to linear mode if it is bombarded with high-intensity light (or, potentially, by heating the APD [69]).
@@
The Bright Illumination attack (shown in [56]) relies on APDs having an undesired, linear mode of operation, and the ability to force an APD to operate in the linear mode by sending high-intensity light to it.
@@
7.2 Is Bright Illumination a Quantum Side-Channel Attack?
@@
7.3 Finding the Vulnerabilities via Quantum Fuzzing
@@
If the Quantum Fuzzing process were to stop after this observation, the useful nature of the linear mode would not be revealed. However, as defined in Section 4.3, if a test case triggers an “interesting” effect on a device, our strategy suggests combining it with other test cases, in the hope of expanding the effect. Employing this strategy, we would send input states that include a high-intensity pulse, followed by another photon pulse. When we test various intensity values for the second pulse, gradually increasing the intensity range and granularity, we reveal two more properties, one after the other:
@@
These two properties reveal how Bob’s device acts when it is “blinded” (that is, when the APDs in Bob’s device are in the linear mode). In fact, these properties reveal Measurement Space Vulnerabilities and Interpretation Vulnerabilities in Bob’s device (as defined in Section 3.2.1): when Bob’s device is blinded, it measures four states (the high-photon-number variants of the BB84 states) that it should not measure, and interprets them as valid if they match his basis choice, or as a loss if they do not match his basis choice.
@@
7.4 Applying Reversed-Space
In the previous subsection, we used Quantum Fuzzing to find a set of Measurement Space Vulnerabilities and Interpretation Vulnerabilities. We now wish to build a Reversed-Space Attack using these vulnerabilities, according to the procedure described in Section 5. For the sake of readability, the computation can be found in Appendix F. The result of the computation is an attack of the form:
@@ 7.4 Applying Reversed-Space
Similarly to Section 5.3, this attack breaks the security (and the robustness: see [61]) of the implementation, because Eve gains full information on the key without inducing any error and without causing the protocol to abort.
@@
8 Applicability to Current Implementation Attacks
@@
To illustrate the relationships between different attacks on QKD implementations and our methodologies, we present a visual diagram in Figure 2, accompanied by an explanation of the results and insights presented in the diagram. High-level details on each attack in the diagram can be found in Appendix G.
Figure 2: A representation of the relations between QKD attacks and our results.
@@
8.1 Relationship to Quantum Side-Channel Attacks
Observation 5.
@@ 8.1 Relationship to Quantum Side-Channel Attacks
However, in contrast to earlier definitions of side-channel attacks in QKD (see, e.g., [55]), many practical attacks are not Quantum Side-Channel Attacks: many (if not most) well-known attack examples do not use elements that are external to Alice and Bob’s realistic protocol, and are therefore not Quantum Side-Channel Attacks, but rather Quantum State-Channel Attacks. Examples of such attacks are the Photon-Number-Splitting attack, the Trojan Pony attack, and the Time-Shift attack. Hence, our definitions of Quantum Side-Channel Attacks and Quantum State-Channel Attacks represent a meaningful distinction between implementation attacks.
8.2 Relationship to Reversed-Space Attacks
Observation 6.
@@ 8.2 Relationship to Reversed-Space Attacks
Reversed-Space Attacks are a generalization of Faked-States attacks for the following reason: since Faked-States attacks rely on the existence of “faked states” that force a specific interpretation in Bob’s measurement device, they rely on Measurement Space Vulnerabilities and Interpretation Vulnerabilities. Hence, all Faked-States attacks are in particular Reversed-Space Attacks. However, the opposite is not true: the Time-Shift attack is a Reversed-Space Attack which is not a Faked-States attack, because it depends on an enlarged measured space, but it does not measure Alice’s signal as is required in Faked-States attacks [19].
@@
8.3 Relationship to Quantum Fuzzing
Observation 7.
@@ 8.3 Relationship to Quantum Fuzzing
9 Conclusions and Future Work
@@
Since our work focuses on attacks on imperfect QKD devices, one may try to partially solve the problem by using QKD protocols with an untrusted center party (typically named Charlie): since an untrusted party is assumed to be malicious, imperfections in its implementation cannot harm security. However, even in such protocols, imperfections in the trusted parties can still harm security, and our tools can still be used to analyze them. While entanglement-based protocols (such as BBM92 [74]) protect against state generation imperfections, attacks against the measurement device (such as Reversed-Space Attacks) are still possible. Similarly, in protocols with measurement by an untrusted center (such as BHM96 [59] or MDI-QKD [5]), attacks against the state preparation devices (such as the Photon-Number-Splitting and Large Pulse attacks [16, 17]) are still possible. While Device-Independent QKD protocols [75] assume all devices to be untrusted and thus theoretically secure against implementation attacks, they still depend on many assumptions, including the assumption that Alice and Bob’s input choices are random and independent. Failure to satisfy these assumptions in practical implementations can, in fact, compromise the security of the protocol, as seen, for example, in [76].
@@ Acknowledgements
The work of R.L. was partly supported by the Government of Spain (Severo Ochoa CEX2019-000910-S, FUNQIP, and European Union NextGenerationEU PRTR-C17.I1), Fundació Cellex, Fundació Mir-Puig, Generalitat de Catalunya (CERCA program), and the European Union (Quantum Security Networks Partnership, Grant 101114043). The work of T.M. was partly supported by the Technion’s Helen Diller Quantum Center (Haifa, Israel).
@@ Appendix A Application of Linear Optical Devices on Photonic States
In this appendix, we discuss physical implementations of QKD schemes using photons. We first describe Fock-space notations, which are easier to use when considering photon-implemented qubits. We then describe the unitary transformation that describes various optical devices commonly used in contemporary QKD implementations, such as beam splitters, phase shifters, and interferometers. We explain the quantum state of a photon (or more generally, a pulse of photons) passing through these optical devices.
@@
A beam splitter has two input arms (modes 1 and 2) and two output arms (modes 3 and 4), see Figure 3. Given that the Beam Slitter is symmetric, Each entering photon has equal amplitude for transmittance and reflection; The transmitted part keeps the same phase as the incoming photon, while the reflected part gets an extra phase of $e^{i\pi/2}$. Specifically, $|{10}\rangle^{\!\text{\tiny F}}_{1,2}\to\frac{1}{\sqrt{2}}(|{10}\rangle^{\!\text{\tiny F}}_{3,4}+i|{01}\rangle^{\!\text{\tiny F}}_{3,4})$ and $|{01}\rangle^{\!\text{\tiny F}}_{1,2}\to\frac{1}{\sqrt{2}}(i|{10}\rangle^{\!\text{\tiny F}}_{3,4}+|{01}\rangle^{\!\text{\tiny F}}_{3,4})$. Thus, for a single photon state in a general superposition between the two input arms, the transformation is given by
$\alpha|{10}\rangle^{\!\text{\tiny F}}_{1,2}+\beta|{01}\rangle^{\!\text{\tiny F}}_{1,2}\mapsto\frac{\alpha+i\beta}{\sqrt{2}}|{10}\rangle^{\!\text{\tiny F}}_{3,4}+\frac{i\alpha+\beta}{\sqrt{2}}|{01}\rangle^{\!\text{\tiny F}}_{3,4}.$ (12)
It is important to note that when a single mode (carrying a single photon) enters a beam splitter from one arm, and nothing (namely, vacuum) enters the other arm (say, $\alpha=1;\beta=0$), there are still two input modes and two output modes. This means that the other (vacuum) entry must be considered as an additional mode — an ancilla carrying no photons.
(1)(3)(2)(4)
Figure 3: A symmetric beam splitter with two input modes (1) and (2) and two output modes (3) and (4).
@@
According to [77], the result of a symmetric beam splitter on a pulse with $n$ photons, denoted as $|{n,0}\rangle_{1,2}$, is given by:
$|{\psi_{\mathrm{out}}}\rangle=\frac{1}{\sqrt{2^{n}}}\sum_{k=0}^{n}\sqrt{\binom{n}{k}}|{k,n-k}\rangle_{3,4}.$ (13)
@@
A controlled phase shifter $P_{\phi}$ performs a phase shift on the input state by a given phase $\phi$, i.e. $P_{\phi}(|{n}\rangle^{\!\text{\tiny F}})=e^{i\cdot n\cdot\phi}|{n}\rangle^{\!\text{\tiny F}}$, see [78].
@@
We will now describe the operation of the interferometer on input states similar to those used in the example protocol in Section 5.3, under the more generic and physically accurate Fock-space notations (Section 2.2).
In each transmission, a superposition of two (time) modes enters the interferometer and result in a superposition of 6 modes (Figure 6). The input modes are separated with a time difference of $\Delta T$ seconds, that is, the first mode arrives at time $t_{0}^{\prime}$, and the second at $t_{1}^{\prime}=t_{0}^{\prime}+\Delta T$. The first pulse travels through the short arm in $T_{\rm short}$ seconds, and through the long arm in $T_{\rm long}=T_{\rm short}+\Delta T$ seconds, where the time difference between the two arms is exactly the time difference $\Delta T$ between the two incoming modes. Due to traveling through both arms, the first mode yields outgoing pulses both at time $t_{0}\equiv t^{\prime}_{0}+T_{short}$ and at $t_{1}\equiv t^{\prime}_{0}+T_{long}=t^{\prime}_{0}+T_{short}+\Delta T=t_{0}+\Delta T$.
When the second pulse enters the interferometer, it also travels through both arms. Intuitively, the part of the $t^{\prime}_{1}$ mode that travels through the short arm interferes with the part of the $t^{\prime}_{0}$ mode that travels through the long arm, and the output exits the interferometer at $t_{1}$. The part of the second pulse that travels through the long arm exits the interferometer at time $t_{2}=t_{1}+\Delta T$. As a result, there are six different possible modes at the two output arms, three in each direction, with the two middle pulses determined by the interference between the two pulses arriving into Bob’s lab. We shall now show the mathematical formulation behind these statements.

### v3

@@ 6.3. Examples
6.3.1. Quantum Side-Channel Attacks
The Large Pulse attack (vakhitov_large_2001, ):
@@ 6.3.1. Quantum Side-Channel Attacks
The Injection-Locking attack (Pang2020, ):
@@
The Bright Illumination attack (lydersen_blinding_commercial_2010, ):
@@
6.3.2. Quantum State-Channel Attacks
The photon-number splitting attack (PNS, ):
@@ 6.3.2. Quantum State-Channel Attacks
The Time-Shift attack (qi_time-shift_2006, ):
@@
Trojan-Pony Attack (gottesman_security_2004, ):
@@
Imperfect Faraday Mirror Attack (wang_faraday_2013, ):
@@
6.4. Application to Security Analyses
As shown in the above subsection, a wide variety of side channels can be used to attack QKD implementations. For an implementation to be proven secure, all these side channels must be considered in the theoretical model of the implementation, though we are not aware of security proofs that use this full modeling. Alternatively, a separate engineering analysis can examine the physical QKD device and its potential side channels and ensure that these side channels cannot be exploited by Eve.
7. Putting It All Together: Analysis of the Bright Illumination Attack
@@
7.1. The Bright Illumination Attack
The “Bright Illumination” attack (makarov_blinding_first_2009, ; lydersen_blinding_commercial_2010, ; lydersen_blinding_thermal_2010, ; sague_blinding_pulsed_2011, ) is an implementation attack that changes the behavior of QKD receivers by exposing them to high-intensity light. This attack was implemented against a variety of QKD implementations, including implementations from recent years (Gras2020, ; Gao2022, ), using various photodetectors and implementing various protocols (lydersen_blinding_snspd_2011, ; chaiwongkhot_blinding_tes_2022, ). In this section we analyze the specific attack and implementation shown in (sague_blinding_pulsed_2011, ); however, our analysis can be applied with minor modifications to other Bright Illumination attacks against various QKD implementations.
@@ 7.1. The Bright Illumination Attack
In the QKD system targeted by (sague_blinding_pulsed_2011, ), the implemented protocol is polarization-based BB84, and measurement is performed using a passive basis choice and four detectors, as shown in Figure 1 (a).
Figure 1. (a) Structure of the QKD receiver targeted in the Bright Illumination attack. PIBS: polarization-independent beam splitter. PBS: polarizing beam splitter. PR: polarization rotator. APD: avalanche photodiode. (b) Propagation of Eve’s high-intensity pulse through the QKD receiver.
(a) shows basis choice by PIBS, and one PBS and two APDs for each basis measurement. In (b), Eve activates only one detector. Subfigure (a) shows a QKD receiver made of a PIBS connected to a PBS on its right and a PR followed by a PBS above. Each PBS is connected to two APDs, one on its right and one above. Subfigure (b) shows the flow of light through the device in a Bright Illumination attack. A wide red arrow, representing light of intensity ”2 P threshold”, enters the PIBS. Two regular-width arrows, representing light of intensity ”P threshold”, then connect the PIBS to each one of the PBSes (the PBS on the right and the PBS above, respectively). For the PBS on the right, the entire arrow (light of intensity ”P threshold”) continues to the right APD, while the APD above does not receive any light, so a check mark appears near the right APD, and an X mark appears near the APD above. In contrast, for the PBS above, the PBS is connected to each APD with a thin-width arrow (each arrow representing light of intensity ”P threshold / 2”), and an X mark appears near each one of those APDs.
@@
APDs have two possible modes of operation, called the “Geiger mode” and “linear mode” (makarov_blinding_first_2009, ). In Geiger mode, an APD operates as a single-photon detector: it generates a “click” (a macroscopic event) when it receives one or more photons. However, in linear mode, the APD “clicks” only if the incoming pulse’s intensity is above a certain threshold: if the pulse intensity is below that threshold, no effect is registered in the APD (there is no click). Thus, an APD in linear mode is “blind” to low-intensity pulses, including single photons.
In QKD setups, APDs operate in Geiger mode as single-photon detectors. However, an APD can be forcibly reconfigured to linear mode if it is bombarded with high-intensity light (or, potentially, by heating the APD (lydersen_blinding_thermal_2010, )).
@@
The Bright Illumination attack (shown in (sague_blinding_pulsed_2011, )) relies on APDs having an undesired, linear mode of operation, and the ability to force an APD to operate in the linear mode by sending high-intensity light to it.
@@
7.2. Is Bright Illumination a Quantum Side-Channel Attack?
@@
7.3. Finding the Vulnerabilities via Quantum Fuzzing
@@
If the Quantum Fuzzing process were to stop after this observation, the useful nature of the linear mode would not be revealed. However, as discussed in Section 4.3, if a test case triggers an “interesting” effect on a device, our strategy suggests combining it with other test cases, in the hope of expanding the effect. Employing this strategy, we would send input states that include a high-intensity pulse, followed by another photon pulse. When we test various intensity values for the second pulse, gradually increasing the intensity range and granularity, we reveal two more properties, one after the other:
@@
These two properties reveal how Bob’s device acts when it is “blinded” (that is, when the APDs in Bob’s device are in the linear mode). In fact, these properties reveal Measurement Space Vulnerabilities and Interpretation Vulnerabilities in Bob’s device (see Definitions 4 and 5): when Bob’s device is blinded, it measures four states (the high-photon-number variants of the BB84 states) that it should not measure, and interprets them as valid if they match his basis choice, or as a loss if they do not match his basis choice.
@@
7.4. Applying Reversed-Space
In the previous subsection, we used Quantum Fuzzing to find a set of Measurement Space Vulnerabilities and Interpretation Vulnerabilities. We now wish to build a Reversed-Space Attack using these vulnerabilities, according to the procedure described in Section 5. For the sake of readability, the computation can be found in Appendix D. The result of the computation is an attack of the form:
@@ 7.4. Applying Reversed-Space
Similarly to Section 5.3, this attack breaks the security (and the robustness: see (BKM2007, )) of the implementation, because Eve gains full information on the key without inducing any error and without causing the protocol to abort.
@@
8. Applicability to Current Implementation Attacks
@@
To illustrate the relationships between different attacks on QKD implementations and our methodologies, we present a visual diagram in Figure 2, accompanied by an explanation of the results and insights presented in the diagram. High-level details on each attack in the diagram can be found in Appendix E.
Figure 2. A representation of the relations between QKD attacks and our results.
Special cases of attacks are displayed at the bottom of the diagram, and are classified into more general classes at the top. The bottom of the diagram consists of nine attacks: Photon-Number Splitting Attack, Injection Locking Attack, Large Pulse Attack: Alice, Large Pulse Attack: Bob, Bright Illumination Attacks, Fixed-Apparatus Attack, Trojan Pony Attack, Time-Shift Attack, and Detector Efficiency Mismatch Attacks. Large Pulse Attacks are a general case of Large Pulse Attack: Alice and Large Pulse Attack: Bob. Trojan Horse Attacks are a general case of Large Pulse Attacks. Faked States Attacks are a general case of Bright Illumination Attacks, Fixed-Apparatus Attack, Trojan Pony Attack, and Detector Efficiency Mismatch Attacks. Reversed-Space Attacks are a general case of Faked States Attacks and Time-Shift Attack. Quantum Side-Channel Attacks are a general case of Injection Locking Attack, Trojan Horse Attacks, and Bright Illumination Attacks. Quantum State-Channel Attacks are a general case of Photon-Number Splitting Attack, Time-Shift Attack, Trojan Pony Attack, and Detector Efficiency Mismatch Attacks. Attacks Against Alice are a general case of Photon-Number Splitting Attack, Injection Locking Attack, and Large Pulse Attack: Alice. Attacks Against Bob are a general case of Large Pulse Attack: Bob and Reversed-Space Attacks. The attacks in the diagram that are novel definitions in the paper are Reversed-Space Attacks, Quantum Side-Channel Attacks, and Quantum State-Channel Attacks. The attacks in the diagram that are discoverable through Quantum Fuzzing are Large Pulse Attack: Bob, Bright Illumination Attacks, Time-Shift Attack, and Detector Efficiency Mismatch Attacks. The Attack Families in the diagram are Bright Illumination Attacks, Detector Efficiency Mismatch Attacks, Large Pulse Attacks, Faked States Attacks, Trojan Horse Attacks, Reversed-Space Attacks, Attacks Against Alice, Attacks Against Bob, Quantum Side-Channel Attacks, and Quantum State-Channel Attacks.
@@
8.1. Relationship to Quantum Side-Channel Attacks
Observation 1.
@@ 8.1. Relationship to Quantum Side-Channel Attacks
However, in contrast to earlier definitions of side-channel attacks in QKD (see, e.g., (Nauerth2009, )), many practical attacks are not Quantum Side-Channel Attacks: many (if not most) well-known attack examples do not use elements that are external to Alice and Bob’s realistic protocol, and are therefore not Quantum Side-Channel Attacks, but rather Quantum State-Channel Attacks. Examples of such attacks are the Photon-Number-Splitting attack, the Trojan Pony attack, and the Time-Shift attack. Hence, our definitions of Quantum Side-Channel Attacks and Quantum State-Channel Attacks represent a meaningful distinction between implementation attacks.
8.2. Relationship to Reversed-Space Attacks
Observation 2.
@@ 8.2. Relationship to Reversed-Space Attacks
Reversed-Space Attacks are a generalization of Faked-States attacks for the following reason: since Faked-States attacks rely on the existence of “faked states” that force a specific interpretation in Bob’s measurement device, they rely on Measurement Space Vulnerabilities and Interpretation Vulnerabilities. Hence, all Faked-States attacks are in particular Reversed-Space Attacks. However, the opposite is not true: the Time-Shift attack is a Reversed-Space Attack which is not a Faked-States attack, because it depends on an enlarged measured space, but it does not measure Alice’s signal as is required in Faked-States attacks (makarov_faked_2005, ).
@@
8.3. Relationship to Quantum Fuzzing
Observation 3.
@@ 8.3. Relationship to Quantum Fuzzing
9. Conclusions and Future Work
@@
Since our work focuses on attacks on imperfect QKD devices, one may try to partially solve the problem by using QKD protocols with an untrusted center party (typically named Charlie): since an untrusted party is assumed to be malicious, imperfections in its implementation cannot harm security. However, even in such protocols, imperfections in the trusted parties can still harm security, and our tools can still be used to analyze them. While entanglement-based protocols (such as BBM92 (BBM92, )) protect against state generation imperfections, attacks against the measurement device (such as Reversed-Space Attacks) are still possible. Similarly, in protocols with measurement by an untrusted center (such as BHM96 (BHM96, ) or MDI-QKD (Xu2020, )), attacks against the state preparation devices (such as the Photon-Number-Splitting and Large Pulse attacks (PNS, ; vakhitov_large_2001, )) are still possible. While Device-Independent QKD protocols (di_qkd, ) assume all devices to be untrusted and thus theoretically secure against implementation attacks, they still depend on many assumptions, including the assumption that Alice and Bob’s input choices are random and independent. Failure to satisfy these assumptions in practical implementations can, in fact, compromise the security of the protocol, as seen, for example, in (Gerhardt2011, ).
@@ Acknowledgements
The work of R.L. was partly supported by the Government of Spain (Severo Ochoa CEX2019-000910-S, FUNQIP, and European Union NextGenerationEU PRTR-C17.I1), Fundació Cellex, Fundació Mir-Puig, Generalitat de Catalunya (CERCA program), and the European Union (Quantum Secure Networks Partnership (QSNP), Grant 101114043). The work of T.M. was partly supported by the Technion’s Helen Diller Quantum Center (Haifa, Israel).
@@ Appendix A Application of Linear Optical Devices on Photonic States
In this appendix, we discuss physical implementations of QKD schemes using photons. We describe the unitary transformations that characterize various optical devices commonly used in contemporary QKD implementations, such as beam splitters, phase shifters, and interferometers. We explain the quantum state of a photon (or more generally, a pulse of photons) passing through these optical devices.
@@
A beam splitter has two input arms (modes 1 and 2) and two output arms (modes 3 and 4), as depicted in Figure 3. Assuming that the beam splitter is symmetric, each entering photon has equal amplitudes for transmittance and reflection; the transmitted part keeps the same phase as the incoming photon, while the reflected part gets an extra phase of $e^{i\pi/2}\equiv i$. Specifically, $|{0{,}1}\rangle^{\!\text{\tiny F}}_{2,1}\to\frac{1}{\sqrt{2}}(|{0{,}1}\rangle^{\!\text{\tiny F}}_{4,3}+i|{1{,}0}\rangle^{\!\text{\tiny F}}_{4,3})$ and $|{1{,}0}\rangle^{\!\text{\tiny F}}_{2,1}\to\frac{1}{\sqrt{2}}(i|{0{,}1}\rangle^{\!\text{\tiny F}}_{4,3}+|{1{,}0}\rangle^{\!\text{\tiny F}}_{4,3})$. Thus, for a single photon state in a general superposition between the two input arms, the transformation is given by
(12) $\alpha|{0{,}1}\rangle^{\!\text{\tiny F}}_{2,1}+\beta|{1{,}0}\rangle^{\!\text{\tiny F}}_{2,1}\mapsto\frac{\alpha+i\beta}{\sqrt{2}}|{0{,}1}\rangle^{\!\text{\tiny F}}_{4,3}+\frac{i\alpha+\beta}{\sqrt{2}}|{1{,}0}\rangle^{\!\text{\tiny F}}_{4,3}.$
It is important to note that when a single photon in a single mode enters a beam splitter from one arm, and nothing — that is, the vacuum state — enters the other arm (say, $\alpha=1;\beta=0$), there are still two input modes and two output modes. This means that the other (vacuum) entry must be considered as an additional mode: an ancilla carrying no photons.
Figure 3. A symmetric beam splitter with two input modes, (1) and (2), and two output modes, (3) and (4).
The Beam Splitter is placed diagonally with numbered inputs and outputs. The edges of the Beam Splitter are at the top-right corner and the bottom-left corner. Input mode (1) is on the left, input mode (2) is at the bottom, output mode (3) is on the right, and output mode (4) is at the top.
@@
As discussed in (Agarwal2012, ), the result of a symmetric beam splitter on a pulse with $n$ photons, denoted as $|{0,n}\rangle^{\!\text{\tiny F}}_{2,1}$, is given by:
(13) $|{\psi_{\mathrm{out}}}\rangle=\frac{1}{\sqrt{2^{n}}}\sum_{k=0}^{n}\sqrt{\binom{n}{k}}|{k,n-k}\rangle^{\!\text{\tiny F}}_{4,3}.$
@@
A controlled phase shifter $P_{\phi}$ performs a phase shift on the input state by a given phase $\phi$ — that is, $P_{\phi}(|{n}\rangle^{\!\text{\tiny F}})=e^{in\phi}|{n}\rangle^{\!\text{\tiny F}}$; see (QOpticsBook, ) for more details.
@@
We will now describe the operation of the interferometer on input states similar to those used in the example protocol in Appendix B.
In each transmission, a superposition of two (time) modes enters the interferometer and result in a superposition of 6 modes (Figure 4). The input modes are separated with a time difference of $\Delta T$ seconds: that is, the first mode arrives at time $t_{0}^{\prime}$, and the second at $t_{1}^{\prime}=t_{0}^{\prime}+\Delta T$. The first pulse travels through the short arm in $T_{\rm short}$ seconds, and through the long arm in $T_{\rm long}=T_{\rm short}+\Delta T$ seconds, where the time difference between the two arms is exactly the time difference $\Delta T$ between the two incoming modes. Due to traveling through both arms, the first mode yields outgoing pulses both at time $t_{0}\equiv t^{\prime}_{0}+T_{\rm short}$ and at $t_{1}\equiv t^{\prime}_{0}+T_{\rm long}=t^{\prime}_{0}+T_{\rm short}+\Delta T=t_{0}+\Delta T$.
When the second pulse enters the interferometer, it also travels through both arms. Intuitively, the part of the $t^{\prime}_{1}$ mode that travels through the short arm interferes with the part of the $t^{\prime}_{0}$ mode that travels through the long arm, and the output exits the interferometer at $t_{1}$. The part of the second pulse that travels through the long arm exits the interferometer at time $t_{2}=t_{1}+\Delta T$. As a result, there are six different possible modes at the two output arms, three in each direction, with the two middle pulses determined by the interference between the two pulses arriving into Bob’s lab.
Figure 4. A Mach-Zehnder interferometer. (a) An input qubit. The time-difference between the two incoming modes is identical to the difference between the two arms; (b) a vacuum state entering the second (blocked) arm; (c) beam splitters; (d) phase shifter $P_{\phi}$; (e) six output modes.
Two possible input pulses, two arms (long and short), and six possible output pulses on the interferometer’s paths. Two pulses denoted ”t prime 0” and ”t prime 1” are at the left input arm of the first beam splitter, denoted (a). A dashed line is at the bottom input arm of the first beam splitter, denoted (b). The short path between the two beam splitters includes nothing, and the long path between them includes a gray circle with a ”phi” symbol denoting a phase shifter (denoted (d)). Both at the right output arm and at the bottom output arm of the second beam splitter, three pulses are illustrated, denoted ”t 2”, ”t 1”, and ”t 0” (a total of six pulses, denoted (e)), where ”t 2” is the closest to the beam splitter, ”t 1” is in the middle, and ”t 0” is the farthest away from the beam splitter. The pulses denoted ”t 1” in both arms are slightly larger than those denoted ”t 0” and ”t 2”.
We shall now show the mathematical formulation behind these statements.